OpenStudy (anonymous):

Okay, I know what to do... but am somehow missing the answer!!! Helppp! 7 (5+0.5x) = 98

OpenStudy (anonymous):

I would multiply the 7 by each term on the left.. 35 + 3.5x = 98

OpenStudy (anonymous):

I tried that... still can't get the answer.

OpenStudy (anonymous):

subtract 35 from both sides and then divide by 3.5 - what did you get?

OpenStudy (anonymous):

3.5x = 98-35 3.5x = 63 x = 63/3.5 x = 18

OpenStudy (anonymous):

Thanx! That's what I was doing. I ALWAYS forget to divide at the end. Thanx so much.

OpenStudy (anonymous):

You are welcome. Happy studying